### v1

# Multi-Scale Cell Decomposition for Path Planning using Restrictive Routing Potential Fields

Abstract
In burgeoning domains, like urban goods distribution, the advent of aerial cargo transportation necessitates the development of routing solutions that prioritize safety. This paper introduces Larp, a novel path planning framework that leverages the concept of restrictive potential fields to forge routes demonstrably safer than those derived from existing methods. The algorithm achieves it by segmenting a potential field into a hierarchy of cells, each with a designated restriction zone determined by obstacle proximity. While the primary impetus behind Larp is to enhance the safety of aerial pathways for cargo-carrying Unmanned Aerial Vehicles (UAVs), its utility extends to a wide array of path planning scenarios. Comparative analyses with both established and contemporary potential field-based methods reveal Larp’s proficiency in maintaining a safe distance from restrictions and its adeptness in circumventing local minima.
I Introduction
As we look to the future, the role of aerial cargo delivery is expected to become central in the domain of urban goods transportation. Given its growing importance, the development of solutions that are both scalable and verifiably safe is crucial [1, 2, 3]. Safety is a paramount concern, and routing solutions are integral to this aspect. The flight paths chosen for unmanned aerial vehicles (UAVs) significantly influence the potential for aerial incidents. Consequently, a route’s function now extends beyond mere connectivity from origin to destination; it must also encompass considerations for safety.
This paper introduces a path planning framework tailored for restrictive routing within potential fields. It decomposes a potential field into multi-scale cells with an estimated maximum potential for their region and utilizes graph path planning algorithms to generate routes that overcome several of the issues with traditional artificial potential field methods. Initially conceived to bolster the safety and scalability of UAV-operated aerial cargo transport within urban air mobility (UAM), the algorithm’s utility extends beyond this scope. It is versatile enough to be implemented in any setting that aligns with the potential field standard to be discussed [4].
I-A Background
A variety of distinct and diverse routing approaches have been proposed for the path planning of UAVs and similar vehicles in complex environments [5]. Two prominent categories are sampling-based and node-based approaches, which involve sampling or deconstructing the environment into networks of nodes or cells for navigation, respectively. Any sector containing an obstacle is excluded from consideration, and the remaining space is searched for optimal routing. Node-based and some sampling-based approaches often employ established graph path planning algorithms to determine routes after decomposing the feasible space. Notable algorithms in these categories include rapidly-exploring random trees (RTTs), artificial potential fields (APFs), A*, and their variants [6, 7, 8, 9, 10, 11, 12].
Another major category gaining popularity is mathematical model-based approaches. These methods formulate and apply mathematical expressions to determine optimal routes either through optimization or by solving numerical equations. A prominent approach in this category is mixed-integer linear programming (MILP) [13, 14].
Bio-inspired algorithms, as the name suggests, leverage intrinsic properties found in nature to design optimal routes. Ant colony optimization (ACO) and particle swarm optimization (PSO) have become standard algorithms within this category [15, 16, 17].
Reinforcement learning-based algorithms are a recent class of algorithms that have emerged from advancements in the field of machine learning. These methods fundamentally optimize a model for sequential decision-making in route search [18, 19, 20]. A subset of these methods has also been trained to work with potential fields for path planning [20]. However, due to the ‘black box‘ nature of data-driven models and their non-deterministic output, machine learning models are not considered for comparison in this research. Routes for cargo UAVs must be reproducible, scalable, and verifiable for practical application.
I-B Contributions
A significant portion of the works reviewed focus on grid-based environments, which may not accurately reflect real-world situations. Moreover, only a handful of these works actively incorporate variable proximity to restrictions as part of their route search. Our work stands out as one of the few to considers the safety of the generated routes and their potential to violate restrictions.
Akin to our work, [12, 21] explore the application of potential fields for path planning, with [12] employing them specifically to quantify safety. These studies also utilize graph-based algorithms such as A* to navigate around local minima, a well-known issue associated with APFs methods. However, the methodology of [12] is confined to grid-based environments, and [21] primarily applies A* within a grid context to overcome local minima, which may lead to routes that are not optimized for distance.
Our work expands upon these concepts by considering both continuous and complex spaces within a unified potential field framework. We examine dynamic proximity to obstacles using multi-resolution cells and assign a safety guideline to routes in both complex and non-complex environments through the concept of potential for restriction violation.
II Preliminaries
II-A Restrictive Routing Potential Fields
In [4], we explore the problem of last-mile UAVs air management and aerial cargo delivery in urban environments, and introduce a standard for restrictive routing based on a repulsive artificial potential field. Assuming a constant operating altitude, physical and virtual restrictions are designated as areas of high potential that diminishes the further an UAV (or agent) is from an obstacle; a behavior dictated by their repulsion matrix.
The standardization facilitates an uniform approach to potential field construction, enabling consistent replication and analysis across different routing algorithms. By leveraging the structured nature of the field, such algorithms can efficiently navigate through complex environments utilizing a common standard, avoiding obstacles, and minimizing the risk of restriction violations. The potential field delineated herein is henceforth denoted as a ‘restrictive routing potential field‘.
II-B Standardized Potential Field Units
Under the proposed standard, a restrictive routing potential field is comprised of multiple building block units, each corresponding to a distinct type of restriction. Upon examining their definitions, a set of intrinsic properties emerges: a repulsion vector $\bar{x}$, a scaled squared distance $\tilde{d}\,^{2}(x)$, and a potential evaluation $\sigma(x)$. The repulsion vector $\bar{x}$ indicates the direction and magnitude of the force repelling from restricted areas. The scaled squared distance $\tilde{d}\,^{2}(x)$ reflects the scaled proximity to a point $x$, modulated by the unit’s repulsion matrix $A$. Finally, the potential $\sigma(x)$ denotes the field’s influence at point $x$, effectively reflecting the potential for restriction violation.
The repulsion vectors for the fundamental units and collection of them are detailed in Table I.
TABLE I: Restrictive Routing Potential Field Units
Unit Parameters Repulsion vector $\bar{x}(x)$ Field
Point
Location $\hat{x}$
Repulsion matrix $\mathrm{A}$
$\bar{x}_{p}(x)=x-\hat{x}$ [Uncaptioned image]
Line
Line start $\hat{x}_{1}$
Line end $\hat{x}_{2}$
Repulsion matrix $\mathrm{A}$
$\begin{matrix}\rho(x)=\frac{(\hat{x}_{2}-\hat{x}_{1})\cdot(x-\hat{x}_{1})}{\|\hat{x}_{2}-\hat{x}_{1}\|^{2}}\\ \\ \bar{x}_{l}(x)=x-\hat{x}_{1}+\mathop{\rm clamp}\left(\rho(x),0,1\right)(\hat{x}_{2}-\hat{x}_{1})\end{matrix}$ [Uncaptioned image]
Rectangle
Corner $\hat{x}_{1}$
Opposite corner $\hat{x}_{2}$
Repulsion matrix $\mathrm{A}$
$\begin{matrix}g(x)=\frac{1}{2}\left(|x-\hat{x}_{1}|+|x-\hat{x}_{2}|-|\hat{x}_{1}-\hat{x}_{2}|\right)\\ \\ \bar{x}_{r}(x)=\mathop{\rm sign}(x-\hat{x}_{1})\odot g(x)\end{matrix}$ [Uncaptioned image]
Ellipse
Location $\hat{x}$
Repulsion matrix $\mathrm{A}$
Ellipse shape matrix $\mathrm{B}$
$\bar{x}_{e}(x)=\mathop{\rm max}\left(1-\frac{1}{\|\mathrm{B}^{-1}\left(x-\hat{x}\right)\|},0\right)\left(x-\hat{x}\right)$ [Uncaptioned image]
Collection Parameters of the sub units Repulsion vector $\bar{x}(x)$ of sub unit with smallest scaled squared distance $\tilde{d}\,^{2}(x)$ [Uncaptioned image]
Definition 1 (Squared Distance to a Potential Field Unit).
A proxy to the squared distance to a potential field unit $d\,^{2}(x)$ is defined as:
$d\,^{2}(x)=\bar{x}(x)^{T}\bar{x}(x),$
where $\bar{x}(x)$ is the repulsion vector detailed in Table I and $x$ is an evaluated point. For a collective of units, $d\,^{2}(x)$ is the smallest squared distance among all sub-units.
Definition 2 (Scaled Squared Distance to a Potential Field Unit).
The scaled squared distance to a potential field unit $\tilde{d}\,^{2}(x)$ is defined as:
$\tilde{d}\,^{2}(x)=\bar{x}(x)^{T}\mathrm{A}^{-1}\bar{x}(x),$
where $\bar{x}(x)$ is again the repulsion vector from Table I, $\mathrm{A}$ is the positive definite repulsion matrix of the unit, and $x$ is an evaluated point. In the case of a collection of units, $\tilde{d}\,^{2}(x)$ is the minimum of the scaled squared distances with respect to its sub-units.
Definition 3 (Potential of an Unit).
The potential of a unit $\sigma(x)$ is defined as the exponential of the scaled squared distance to the unit:
$\sigma(x)=\exp{\left(\tilde{d}\,^{2}(x)\right)},$
where $\tilde{d}\,^{2}(x)$ is the scaled squared distance, and $x$ is an evaluated point.
III Methodology
Given the properties of the standardized potential field, we introduce Larp (Last-mile restrictive path planning) as a routing framework for safe path planning. It consists of three core stages: first, an primary algorithm that decomposes a field into distinct multi-scale cells; secondly, the formation of a routing network using the cells and an adjacency detection algorithm; lastly, it traverses the graph to determine a suitable route. The stages can be observed in Fig. 1. An official implementation of Larp can also be found on our Github repository at https://github.com/wzjoriv/Larp.
(a) Potential field
(b) Quad cells
(c) Routing network
(d) A route in the network
Figure 1: Stages of Larp for a walled room scene
III-A Multi-Scale Cell Decomposition
Drawing inspiration from adaptive quad tree cell decomposition, the first phase consists of partitioning a potential field into a multitude of cells, each varying in size. These cells are then correlated with their respective potential for infringing upon nearby restrictions. The stratagem for cell subdivision is calibrated, ensuring that cells proximal to obstacles are proportionately smaller, thereby augmenting routing precision and safety. Fig. 1(b) illustrates this decomposition for a walled room scene. The methodology behind cell decomposition is expounded in Algorithms 1 and 2.
Algorithm 1 Larp – Quad Tree Cell Decomposition
minimum cell size $n_{min}$; maximum cell size $n_{max}$
Cells containing restriction have highest potential
function Build($x$, $n$, $U$)
     $quad\leftarrow QuadNode(x,n)$
     $zones\leftarrow$ map with default farthest zone
     if $|U|>0$ then
         $zones\leftarrow$ ApproxDistanceZones($x$, $n$, $U$)
     end if
     $quad.zone\leftarrow\min(zones)$triangleright Farthest zone if empty
     if $n\leq n_{max}$ then
         if $n\leq n_{min}$ or $quad.zone$ is farthest zone then
              Mark $quad$ as leaf
              return $quad$
         else if $quad.zone>0$ then
[…]
              $\bar{C}\leftarrow\{c=x+\frac{n}{\sqrt{2}\,|\bar{x}_{u}(x)|}\,\bar{x}_{u}(x)\,|\,u\in\,\bar{U}\}$
              $bd_{l}\leftarrow boundary\_egde[quad.zone]$
              for each point $c$ in $\bar{C}$ do
                  if given any $u\in\bar{U}$, $\tilde{d}_{u}^{2}(x)<bd_{l}$ then
                       Mark $quad$ as leaf
                       return $quad$
                  end if
              end for
         end if
     end if
[…]
     for each quad $q$ of a cell do
         $quad.child[q]\leftarrow$ Build($q.center$, $n/2$, $U$)
     end for
     return $quad$
end function
$root\leftarrow$ BUILD(field.center, field.size, all units in field)
Algorithm 2 Larp – Estimate Boundary of Cell
Sorted list of boundaries $bds$ in ascending order
function ApproxDistanceZones($x$, $n$, $U$)
     $zones\leftarrow$ map with default further zone
     for each unit $u$ in $U$ do
         if $d_{u}^{2}(x)\leq\frac{n^{2}}{2}$ then
              $zones[u]\leftarrow 0$
         end if
     end for
     for each unit $u$ in $U$ but not already in $zones$ do
         $dist=\tilde{d}_{u}^{2}(x)$
         Assign $zones[u]$ into bin given $bds$ and $dist$
     end for
     return $zones$
end function
III-B Routing network
To streamline path planning within the designated space, a network graph is constructed subsequent to the potential field’s decomposition. The process employs a divide-and-conquer approach to traverse the quad tree, facilitating the identification of neighboring nodes for any given leaf quad node. The intricacies of the field are distilled into a simplified network, where each graph node correlate with a leaf cell and its upper bound potential for the vicinity. Fig. 1(c) provides a visual representation of this network graph.
In the interest of brevity, we have omitted the detailed adjacency algorithm used to construct the routing network; however, it is available for review in our public repository.
III-C Path Planning
Upon the establishment of the network, the implementation of any graph routing algorithm that simultaneously minimizes route distance and accumulated potential is feasible. For this study, we employ a modified variant of the renowned A* algorithm, wherein the distance function is redefined as:
$d(q_{a},q_{b})=s(q_{b})\cdot\|q_{a}.\text{center}-q_{b}.\text{center}\|$ (1)
where $q_{a}$ and $q_{b}$ denote the quadrant being traversed from and to, respectively; $s(q)$ represents a hyper-parameter scale transformation function that modulates the safety distance; and $q_{a}.\text{center}$ signifies the centroid of cell $q_{a}$. A route generated for the walled room scene is depicted in Fig. 1(d). Subsequent to determining a routing solution, line segments are integrated to connect the initial location and the destination with the discovered route within the network.
III-D Safety Validation
Prior to assessing the safety of a route, it is imperative to define what constitutes a route within a potential field.
Definition 4 (Route in a Potential Field).
A route $R$ within a restrictive routing potential field is delineated as a trajectory comprising a line string that interconnects a sequence of points, expressed as:
$R_{p}=\{x_{1},x_{2},...,x_{N}\}$
where $x_{i}$ signifies a 2D coordinate in a plane, and $N-1$ represents the count of consecutive line segments that constitute the line string.
With Definition 4 established, several metrics to evaluate the safety of a route can be examined. The primary metric is the cumulative potential under a route’s trajectory, calculated as:
[…]
Conversely, a route’s length as a measure of safety is considered:
[…]
The extended duration of a route implies a prolonged presence within the field, potentially elevating the risk of violation occurrences.
Ultimately, combining these two metrics yields a composite measure that reflects the average engagement with restrictions along a route
$R_{avg}=\frac{R_{A}}{R_{d}}$ (4)
A heightened $R_{avg}$ value suggests a route’s active proximity to restricted areas. Empirical observations indicate that a threshold of 0.35 or below is typically indicative of minimal engagement with restrictions.
IV Experiments and Results
IV-A Potential field-based path planning algorithms
To assess the efficacy of our algorithm, we have compiled a suite of path planning strategies that leverage potential fields to evaluate proximity to obstacles and violating routing restrictions. Detailed implementations and comparisons of these methods are available in our Github repository at https://github.com/wzjoriv/path-planning-pf.
IV-A1 Penalty Method with Gradient Descent (PM)
The penalty method with gradient descent introduces an attractive force that propels an agent towards a goal, defined as:
$F_{A}(x)=\zeta(x_{g}-x)$ (5)
where $\zeta$ is a scalar hyper-parameter dictating the strength of the attraction, and $x_{g}$ denotes the goal’s location. Additionally, a repulsive force emerges from a penalty on the potential field value,
$F_{R}(x)=-\eta\,p^{\prime}(\sigma(x))\,\nabla\sigma(x)$ (6)
where $\eta$ is a scalar hyper-parameter governing the repulsion, $p^{\prime}(\cdot)$ is the derivative of the penalty function, and $\nabla\sigma(x)$ represents the gradient of the potential field under discussion.
Combining these two forces forms a traversal strategy that steers the agent towards the goal, articulated as:
$x_{i+1}=x_{i}+\gamma\,\frac{F_{A}(x_{i})+F_{B}(x_{i})}{\|F_{A}(x_{i})+F_{B}(x_{i})\|}$ (7)
where $\gamma$ signifies the step size, and $x_{0}$ is the agent’s initial location. To circumvent local minima near the goal, our experiments introduce a heuristic: if $\|x_{N}-x_{g}\|\leq 2.5$, a direct line segment to the goal $x_{g}$ is appended to the route’s $R_{p}$.
IV-A2 Artificial Potential Field (APF, APF(*))
The classical APF method supplants the gradient-based approach with a traditional repulsive force as delineated in [10, 22]. Within the scope of the problems addressed, this repulsive force is articulated by:
$F_{R}(x)=\begin{cases}\eta\left(\frac{1}{d(x)}-\frac{1}{d_{o}}\right)\frac{1}{d\,^{2}(x)}\bar{x}(x),&\text{if }d(x)\leq d_{o}\\ 0,&\text{otherwise}\end{cases}$ (8)
where $d(x)=\sqrt{d\,^{2}(x)}$, $d_{o}$ is the threshold distance at which the repulsive force becomes active, and $\bar{x}(x)$ is the repulsion vector from the nearest obstacle.
Subsequent enhancements to the APF methodology, not originally included but recommended for improved efficacy, propose an alternative attraction force [22]:
$F_{A}(x)=\begin{cases}\zeta(x_{g}-x),&\text{if }\|x-x_{g}\|\leq d_{g}\\ \zeta\,d_{g}\frac{x_{g}-x}{\|x-x_{g}\|},&\text{otherwise}\end{cases}$ (9)
where $d_{g}$ is the critical distance beyond which the attraction force adopts a quadratic profile to ensure a smooth convergence to the target.
For the APF method, a variant denoted as APF (*) is also implemented for which square distance function $\mathbf{\text{d}}\,^{2}(x)$ is substituted with its scaled counterpart $\tilde{d}\,^{2}(x)$.
IV-A3 Modified Artificial Potential Field (M-APF)
M-APF, an advanced iteration of APF introduced by [8, 23], innovates upon the conventional repulsion force to mitigate the issue of local minima. The reformulated force is expressed by:
$\begin{split}F_{R1}(x)&=\eta\left(\frac{1}{d(x)}-\frac{1}{d_{o}}\right)\|x-x_{g}\|\,^{m-3}\\ F_{R2}(x)&=\eta\,m\left(\frac{1}{d(x)}-\frac{1}{d_{o}}\right)^{2}\|x-x_{g}\|\,^{m}\\ \\ F_{R}(x)&=\begin{cases}F_{R1}(x)+F_{R2}(x)\frac{\bar{x}(x)}{\|\bar{x}(x)\|},&\text{if }d(x)\leq d_{o}\\ 0,&\text{otherwise}\end{cases}\end{split}$ (10)
where $m$ is a positive scalar hyper-parameter that modulates the repulsion force’s sensitivity to the agent’s proximity to the goal.
IV-B Results
In evaluating these algorithms, we explored three distinct scenarios of path planning.
IV-B1 Scenario 1 (Unobstructed Path)
The first scenario considers a straightforward path planning task where the objective is to travel towards an unobstructed target. Although there are no significant obstacles present, this scenario serves to accentuate the subtle differences between the algorithms. As shown in Table II and Fig. 2, even in an ideal situation where no obstacles impede the path, Larp—placing a premium on safety—may yield routes that are marginally longer if it results in a lower overall potential energy.
TABLE II: Performance metrics for Scenario 1. All algorithms successfully reach the goal. Larp generates a slightly longer path due to its conservative movements and emphasis on safety. A $\checkmark$ under ’Goal Found’ indicates $\|x_{N}-x_{g}\|^{2}\leq 2.5$ where $x_{N}$ is the final position of a route.
Algorithm Goal Found Route Distance Route Area Average Potential Highest Potential
PM 62.04 0.4006 0.0065 0.0619
APF 62.06 0.3863 0.0062 0.0594
APF (*) 62.02 0.4020 0.0065 0.0622
M-APF 62.02 0.4020 0.0065 0.0622
Larp (Our) 62.9181 0.1082 0.0017 0.0084
(a) PM
(b) APF
(c) APF (*)
(d) M-APF
(e) Larp (Our)
Figure 2: Comparison of routes in Scenario 1, featuring an unobstructed path. The red marker indicates the initial location of the agent.
IV-B2 Scenario 2 (Obstructed Path and Near-Obstacle Goal)
This scenario examines the algorithms’ performance when an agent must navigate around an obstacle and when the target is in close proximity to one. The results are presented in Table III and Fig. 3. When obstacles are introduced along the route, both PM and APF struggle to maintain a safe distance from the constraints, resulting in higher peak potential values. PM’s dependence on gradients and penalties hinders its ability to significantly deviate from the obstacle’s path. APF’s higher potential can be attributed to its focus on immediate obstacle proximity rather than the overall potential for obstacle avoidance; however, it also contributes to its shorter route distance. Among all the methods evaluated, Larp consistently produced safer routes, albeit at the expense of increased distance. Nevertheless, Larp’s route lengths remain competitive, particularly when compared to methods such as APF (*) and M-APF.
TABLE III: Performance metrics for Scenario 2. The highest potential energy values for APF (*), M-APF, and Larp reflect the goal’s proximity to obstacles.
Algorithm Goal Found Route Distance Route Area Average Potential Highest Potential
PM 79.188 23.1432 0.2923 0.9011
APF 77.2863 23.011 0.2977 0.9220
APF (*) 87.8328 29.3081 0.3337 0.8948
M-APF 80.8579 21.0194 0.26 0.8948
Larp (Our) 83.8675 15.8687 0.1892 0.8948
(a) PM
(b) APF
(c) APF (*)
(d) M-APF
(e) Larp (Our)
Figure 3: Scenario 2 of an obstructed path and near-obstacle goal (routes comparison)
IV-B3 Scenario 3 (Walled-Off Path)
The final scenario investigates the behavior of the methods when the direct route to a target is obstructed by a wall. As depicted in Table IV and Fig. 4, all methods, except for Larp, fail to reach the destination. When confronted with non-circular restrictions, both penalty-based and force-based methods may struggle to overcome the local minima created by the opposing forces of attraction and repulsion from the obstacles.
TABLE IV: Performance metrics for Scenario 3. Except for Larp, all other methods are unable to reach the target; therefore, their performance measurements are not valid for comparison but provide insight into their behavior. In this table, an ✗  under ’Goal Found’ indicates $\|x_{N}-x_{g}\|^{2}>2.5$
. Algorithm Goal Found Route Distance Route Area Average Potential Highest Potential PM 33.32 18.6015 0.5583 0.9529 APF 42.83 26.3972 0.6163 0.9075 APF (*) 80.37 53.9457 0.6712 0.8075 M-APF 38.57 1.8687 0.0485 0.0674 Larp (Our) 163.8382 14.8769 0.0908 0.3453
(a) PM
(b) APF
(c) APF (*)
(d) M-APF
(e) Larp (Our)
Figure 4: Routes comparison for Scenario 3, featuring a walled-off path.
V Conclusion
In this work, we have presented Larp, an innovative path planning framework tailored for restrictive routing within potential fields. Demonstrating superior results, Larp outperforms both traditional and contemporary potential field-based methodologies by reliably crafting safer routes that diminish the risk of restriction violations, all while preserving comparative efficiency in terms of travel distance. This is achieved through an initial segmentation of the potential field into multi-scale cells, with each cell’s restriction zone determined by its proximity to nearby obstacles. For transparency and to facilitate further research, the source code for an official implementation of Larp has been made publicly accessible.
While Larp marks an advancement in route safety, opportunities for improvement remain. The framework’s quad cell decomposition approach and zone classification system are poised for further optimization. The influence of a cell’s dimensions and positioning on the placement of subsequent cells is a critical factor in the pursuit of optimally short and safe routes. Moreover, the current method of zone classification, which relies on an proxy scaled distance to a cell’s periphery to gauge the highest potential across the region, may estimate an higher potential for a region that what it is. Future endeavors will concentrate on refining these elements and developing a dynamic and scalable system capable of storing and updating cells in real-time to adapt to evolving restrictions for unmanned aircraft system traffic management (UTM).